OpenStudy (anonymous):

confused on this one because if i multiply the first equation by 2, it cancels out both the x and y components. Solve by the elimination method. 2x + 3y =1 4x + 6y = 2

OpenStudy (anonymous):

that is okay but you need to multiply first equation by -2

OpenStudy (anonymous):

sorry about that...you will get 0=0 which is always true so infinitely many or all real numbers as solution

OpenStudy (anonymous):

ok, that is the first time i saw the answer come out like that, thanks!

OpenStudy (anonymous):

you're welcome....good job on elimination...you may also get something like -2=0 which is never true so NO SOLUTION

OpenStudy (anonymous):

so which would it be, infinitely many solutions or no solution?

OpenStudy (anonymous):

no solution is a different type-when left side doesn't equal right side INFINETLY MANY/ALL real numbers mean same things

OpenStudy (radar):

You have what is known as an identity.

OpenStudy (anonymous):

exactly radar

OpenStudy (radar):

I hope the course that ms elegant prepared for such an event lol

OpenStudy (anonymous):

can you help me with polynomial question Find polynomial P(n)=Sn for 1, 2, 3, ...10 and explain method

OpenStudy (anonymous):

so the answer i get when multiplying by -2 is 0=-2. so I'm guessing that the answer is there is no solution. but how do i perform a check to make sure that my answer is right?

OpenStudy (radar):

It is unsolvable,it is an identity, it is the same equation just one is double the other you need two distinct equations for the two unknowns.

OpenStudy (radar):

btm2011, I wouldn't be much help on that polymonial series!

OpenStudy (anonymous):

ok, so there is no solution, thx!

OpenStudy (radar):

Ms elegant do you understand even though they provided you with two equations, they are actually the same, one does not provide any new information for you.

OpenStudy (anonymous):

for that problem the answer is all real numbers, infinitely many (identity) because 0=0 -2(2x+3y=1)____-4y-6y=-2

OpenStudy (anonymous):

-4y-6y=-2 + 4x+6y=2

OpenStudy (anonymous):

There is a solution!

OpenStudy (anonymous):

ok i think i got it, there is a solution, but it is just an infinite amount of solutions that can be the answer, right?

OpenStudy (anonymous):

YES...bc every number equals itself

OpenStudy (radar):

yes, x=0 and y=1/3 could be a solution, but there is no specific solution, I could put down another solution like x=-2, y=1

OpenStudy (anonymous):

ok, i get it. and that's the check, that several answers fit this equation - so there are infinite solutions, not just one

OpenStudy (radar):

That is what it seems to be.

OpenStudy (anonymous):

ok great thx! i appreciate the explanation so that i could understand what i was doing

OpenStudy (anonymous):

GREAT!